Suppose that Ramos contributes $5000/year into a traditional IRA earning interest at the rate of 2%/year compounded annually, every year after age 37 until his retirement at age 67. At the same time, his wife Vanessa deposits $3500/year (the amount after paying taxes at the rate of 30%) into a Roth IRA earning interest at the same rate as that of Ramos. Suppose that Ramos withdraws his investment upon retirement at age 67 and that his investment is then taxed at 30%. (Round your answers to the nearest cent.)

(a)

How much will Ramos's investment be worth (after taxes) at that time?

$

(b)

How much will Vanessa's investment be worth at that time?

$
Answer #1

Considering the tax bracket during contribution years and withdrawn year at the same level, investment amount for Ramos will be similar to that of Vanessa.

For Ramos:

Interest rate is 2% compounding annually.

Example: 5000*0.02 = 100, total investment at the end of year 1 = 5000+100=5100

In year 2 additional 5000 will be invested, hence 5,100+5000= 10,100,

Interest for year 2 would be 10,000*0.02= 202, total investment at the end of year 2 = 10,100+202=10,302 and so on.

For Vanessa:

Since as per Roth IRA, 30% tax was charged on Vanessa's contribution of $ 5000/year, hence total investment per year would be $ 3,500.

3,500*0.02 = 70, total investment at the end of year 1 = 3,500+70=3,570

In year 2 additional 3,500 will be invested, hence 3,570+3,500= 7,070

Interest for year 2 would be 7,070*0.02= 141.4, total investment at the end of year 2 = 7,070+141.4=7,211.4 and so on.

The calculations are shown in the table below:

Ramos contribution Traditional IRA Vanessa contribution ROTH IRA
Year Investment Interest O/s at the end of year Year Investment Interest O/s at the end of year
1          5,000.00             100.00                             5,100.00          1        3,500.00          70.00        3,570.00
2       10,100.00             202.00                           10,302.00          2        7,070.00       141.40        7,211.40
3       15,302.00             306.04                           15,608.04          3      10,711.40       214.23      10,925.63
4       20,608.04             412.16                           21,020.20          4      14,425.63       288.51      14,714.14
5       26,020.20             520.40                           26,540.60          5      18,214.14       364.28      18,578.42
6       31,540.60             630.81                           32,171.42          6      22,078.42       441.57      22,519.99
7       37,171.42             743.43                           37,914.85          7      26,019.99       520.40      26,540.39
8       42,914.85             858.30                           43,773.14          8      30,040.39       600.81      30,641.20
9       48,773.14             975.46                           49,748.60          9      34,141.20       682.82      34,824.02
10       54,748.60         1,094.97                           55,843.58        10      38,324.02       766.48      39,090.50
11       60,843.58         1,216.87                           62,060.45        11      42,590.50       851.81      43,442.31
12       67,060.45         1,341.21                           68,401.66        12      46,942.31       938.85      47,881.16
13       73,401.66         1,468.03                           74,869.69        13      51,381.16    1,027.62      52,408.78
14       79,869.69         1,597.39                           81,467.08        14      55,908.78    1,118.18      57,026.96
15       86,467.08         1,729.34                           88,196.43        15      60,526.96    1,210.54      61,737.50
16       93,196.43         1,863.93                           95,060.35        16      65,237.50    1,304.75      66,542.25
17     100,060.35         2,001.21                        102,061.56        17      70,042.25    1,400.84      71,443.09
18     107,061.56         2,141.23                        109,202.79        18      74,943.09    1,498.86      76,441.96
19     114,202.79         2,284.06                        116,486.85        19      79,941.96    1,598.84      81,540.79
20     121,486.85         2,429.74                        123,916.59        20      85,040.79    1,700.82      86,741.61
21     128,916.59         2,578.33                        131,494.92        21      90,241.61    1,804.83      92,046.44
22     136,494.92         2,729.90                        139,224.82        22      95,546.44    1,910.93      97,457.37
23     144,224.82         2,884.50                        147,109.31        23    100,957.37    2,019.15    102,976.52
24     152,109.31         3,042.19                        155,151.50        24    106,476.52    2,129.53    108,606.05
25     160,151.50         3,203.03                        163,354.53        25    112,106.05    2,242.12    114,348.17
26     168,354.53         3,367.09                        171,721.62        26    117,848.17    2,356.96    120,205.13
27     176,721.62         3,534.43                        180,256.05        27    123,705.13    2,474.10    126,179.24
28     185,256.05         3,705.12                        188,961.17        28    129,679.24    2,593.58    132,272.82
29     193,961.17